Find the degree measure of ∠B.
Given m∠B = 2x+46, m∠E = 3x-1,
ΔABE is a right triangle with m∠A = 90°
m∠B = ___°

Answer by Edwin McCravy(20054)


Since this is a right triangle, the sum of the
measures of the two acute angles is 90°.
m∠B + m∠E = 90°
2x+46 + 3x-1 = 90
5x+45 = 90
5x = 45
x = 9
m∠B = 2x+46 = 2(9)+46 = 18+46 = 64°
Edwin